OpenStudy (anonymous):

Is the simplified form of 2square root of 3 • square root of 12 rational?

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

2square root of 3 * square root of 12 = 2(square root of 3 )* square root of (2*2*3) = 2(square root of 3 )*2* square root of (3) = 2*2 * square root of (3) *square root of (3) =2*2* 3 =12 now can you judge whether it is rational or irrational

OpenStudy (anonymous):

rational right

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

yup correct

OpenStudy (anonymous):

yes!